OpenStudy (anonymous):

A line passes through the point 6,-8 and has a slope of 3/2. Write an equation in slope-intercept form for this line.

OpenStudy (cybershadow):

y = mx + c is the slope-intercept form where m is slope and c is y-intercept and and y and x are x and y coordinates respectively,u can fill in the values in this equation that will give the correct answer u can ask if ya have any doubts @atlantalast

OpenStudy (anonymous):

is it gonna be -8=3/2+6?

OpenStudy (anonymous):

@cybershadow

OpenStudy (anonymous):

yes thats the answer

OpenStudy (cybershadow):

see the equation is y = mx +c c = y-intercept we dont know its value yet but we can find its value -8 = (3/2)6 +c c =-8 -9 c =-17 now put this value in y = mx +c equation that will make it y = (3/2)x -17 thats the answer